# Time Entangled Quantum Blockchain with Phase Encoding for Classical Data

### v1

@@ 2 Time Temporal GHZ State Blockchain
$|B_{xy}\rangle=\frac{1}{\sqrt{2}}(|0\rangle|y\rangle+(-1)^{x}|1\rangle|\tilde{y}\rangle)$
@@ 3 Quantum Blockchain on a Weighted Hypergraph
This scheme was presented in [10] by Shreya Banerjee, A. Mukherjee, and P. K. Panigrahi. Quantum hypergraphs [20] are a group of highly entangled multiparty quantum states. Vertices are quantum particles where edges represent entanglement. Edges encompass more than two vertices and are known as "hypergraphs." With a quantum hypergraph with $k$-hyperedge (a hyperedge connecting $k$ qubits) and $n-1$ vertices, a corresponding quantum state can be prepared [20]. To add the $n$th qubit to the state, initialize the qubit to the Hadamard state: $(|0\rangle+|1\rangle)/\sqrt{2}$. A controlled $Z$ operation is performed with the existing $n-1$ qubits as control and $n$ as the target. For a mathematical hypergraph with five vertices, A, B, C, D, E, a 3-hyperedge over vertices A, B, and C; and a 5-hyperedge over all five vertices, the state is expressed as following.
@@
[21] describes how local unitary operations carried through classical communication (LOCC) does not alter the entanglement of the state under consideration. It is shown in [21] that local applications of unitary Pauli operations on the $n$th qubit one can remove all the $(N-1)$ edges for the special case where an n-hypergraph ($n$-qubit) contains only an $n$ hyperedge. As presented in [22, 23], a weighted hypergraph is when hyperedge carries weights. [22, 24] introduced weighted quantum hypergraphs as locally maximally entangleable (LME) states. In the notation below, $|x\rangle$ is the computational basis, and $f(x)$ is a real number.
@@
In the proposed blockchain, the classical ledger and cryptographic functions have been replaced by the weighted quantum hypergraph and entanglement. They have used the phases (weights) on the hyperedges to encode classical information. Each classical block is a string $P_{i}$ ($i$th classical block) of bits. The block creator initializes the qubit (one qubit represents one quantum block) to $(|0\rangle+|1\rangle)/\sqrt{2}$. Then via a secret bijective function chosen known only by him, $P_{i}$ is converted into the phase $\theta_{P_{i}}$, and the following rotation is applied to the qubit to obtain the $i$th quantum block.
@@
where $n\in\mathbb{N}\setminus\{1\}$. The series converges to $\frac{n}{n-1}$. Consensus can therefore be defined with any such series with an appropriate $\theta_{P_{1}}$. After the consensus execution, peers add the $n$th block to their local chain of $n-1$ blocks. A $C^{(n-1)}Z$ is applied as before with the new block as the target to create the new $n$-hyperedge and a local Pauli-$X$ on the new block to remove the previous $n-1$-hyperedge.
@@ 4.2 Blockchain Protocol
In the new blockchain, for the classical block i, i.e. $P_{i}$, phase $\theta_{P_{i}}$ is calculated by the block creates much the same as in the hypergraph blockchain via a secret bijective function known only to the block creator. However, the outcome of this function is this phase $+$ a classical $2$-bit string $r_{1}r_{2}$. There are two qubits per block, initially in the ground state $|00\rangle$. The following state is then obtained, i.e., the corresponding Bell state for the classical bit string.
@@ 4.4 Consensus
To check the validity of this block, each receiving node measures the quantum block$(\frac{|0r_{m_{1}}\rangle+(-1)^{r_{m_{1}}}e^{i\theta_{m}}|1\bar{r}_{m_{1}}\rangle}{\sqrt{2}})$ in the basis $\frac{|0r_{m_{1}}\rangle\pm(-1)^{r_{m_{1}}}e^{i\theta_{m}\__{pre}}|1\bar{r}_{m_{1}}\rangle}{\sqrt{2}}$ + the other 2 corresponding 2 states that together with these 2 states form an orthonormal basis. If the outcome is not deterministic with the probability of $1$, the phase is invalid, and the block creator is deemed untrustworthy, and the node shares his judgment with the network along with measurement results: the phase is invalid, different nodes have received different phases/states, or different nodes have received different classical strings. With the knowledge of the bit string and the predetermined phase based on $m$, $\theta_{1}$, and ratio $n$, a node can reconstruct the measurement basis for a number of measurements to calculate this probability. In this basis, $\theta_{m}\__{pre}=\frac{\theta_{1}}{n^{(m-1)}}$; $\theta_{P_{1}}$ is shared openly by the creator of the genesis block. If valid, the last copy can be appended to the local blockchain.
@@ 4.5.1 The Spatial GHZ Blockchain Security: Without the Temporal Aspect
The data is encoded into the phase of the GHZ state and the basis states themselves. Therefore, for an adversary to measure the phase, they will need to keep measuring the state in random basis as long as the length of the chain, $\theta_{P_{1}}$, classical 2-bit strings, or $n$ remain secrets. This randomness includes GHZ basis states computational basis states as well. Even if the chain is constructed by the adversary using those parameters (even if he obtains bit strings and the phases of blocks), without the knowledge of the bijective functions, it is not possible to obtain the data. Even if the data is obtained, i.e, the bijective functions are discovered, it is still impossible to mutate the data with that same bijective function intact, i.e, change the phase of the $i$th block - $\theta_{Pi}$ and the classical string - $r_{i_{1}}r_{i_{2}}$, since this would result in a different phase than that which is predefined for the GHZ state and different basis states that don’t represent the classical bit strings $r_{i_{1}}r_{i_{2}}$. If the length of the chain is $m$, i.e. $2m$ qubits, the projection measurement basis to check validity will be
@@
with the rest of the basis set: the other $22^{2n}-$ 2 states that make an orthonormal basis state with these 2 states, where in the projection $\Omega,I-\Omega$}, $\Omega$ is the projection operator for the $+$ state out of the two states above.

### v3

@@ 1 Introduction
  • Absolute Security / Unconditional Security: Security that cannot be compromised using unlimited computational capabilities. Classical security is conditional on the limitations of computational capabilities: dependent on the hardness of a computationally intractable mathematical problem. Information-theoretic security is the idea that needs to be propagated.
  • Temporal GHZ State: A GHZ state where particles are entangled not only in space but also in time.
  • Phase Encoding: Mapping the classical block into a phase angle and adding that phase to Bell and GHZ states.
  • Efficiency and Scalability:: the time-entangled GHZ blockchain encodes a classical block of 2 bits to a quantum block of 2 qubits. Ours is more efficient in the sense that a classical block of arbitrary size in encoded to a quantum block of 2 qubits. Therefore, due to this efficiency, we consider this protocol more scalable since it yields far less quantum information than the original GHZ blockchain [1] for a given amount of classical data. However, not as efficient as the quantum hypergraph blockchain [2], which encodes a classical block of arbitrary size to a quantum block of 1 qubit.
@@ 2 Time Temporal GHZ State Blockchain
[…]
@@ 3 Quantum Blockchain on a Weighted Hypergraph
This scheme was presented in [2] by Shreya Banerjee, A. Mukherjee, and P. K. Panigrahi. Quantum hypergraphs [20] are a group of highly entangled multiparty quantum states. Vertices are quantum particles where edges represent entanglement. Edges encompass more than two vertices and are known as "hypergraphs." With a quantum hypergraph with $k$-hyperedge (a hyperedge connecting $k$ qubits) and n1n-1 vertices, a corresponding quantum state can be prepared [20]. To add the $n$th qubit to the state, initialize the qubit to the Hadamard state: $(|0\rangle+|1\rangle)/\sqrt{2}$. A controlled $Z$ operation is performed with the existing n1n-1 qubits as control and $n$ as the target. For a mathematical hypergraph with five vertices, A, B, C, D, E, a 3-hyperedge over vertices A, B, and C; and a 5-hyperedge over all five vertices, the state is expressed as following.
@@
[21] describes how local unitary operations carried through classical communication (LOCC) does not alter the entanglement of the state under consideration. It is shown in [21] that local applications of unitary Pauli operations on the $n$th qubit one can remove all the (N1)(N-1) edges for the special case where an n-hypergraph ($n$-qubit) contains only an $n$ hyperedge. As presented in [22, 23], a weighted hypergraph is when hyperedge carries weights. [22, 24] introduced weighted quantum hypergraphs as locally maximally entangleable (LME) states. In the notation below, $|x\rangle$ is the computational basis, and $f(x)$ is a real number.
@@
In the proposed blockchain, the classical ledger and cryptographic functions have been replaced by the weighted quantum hypergraph and entanglement. They have used the phases (weights) on the hyperedges to encode classical information. Each classical block is a string $P_{i}$ (iith classical block) of bits. The block creator initializes the qubit (one qubit represents one quantum block) to $(|0\rangle+|1\rangle)/\sqrt{2}$. Then via a secret bijective function chosen known only by him, $P_{i}$ is converted into the phase $\theta_{P_{i}}$, and the following rotation is applied to the qubit to obtain the $i$th quantum block.
@@
where $n\in\mathbb{N}\setminus\{1\}$. The series converges to $\frac{n}{n-1}$. Consensus can therefore be defined with any such series with an appropriate $\theta_{P_{1}}$. After the consensus execution, peers add the $n$th block to their local chain of n1n-1 blocks. A $C^{(n-1)}Z$ is applied as before with the new block as the target to create the new $n$-hyperedge and a local Pauli-$X$ on the new block to remove the previous n1n-1-hyperedge.
@@ 4.2 Blockchain Protocol
In the new blockchain, for the classical block i, i.e. $P_{i}$, phase $\theta_{P_{i}}$ is calculated by the block creates much the same as in the hypergraph blockchain via a secret bijective function known only to the block creator. However, the outcome of this function is this phase ++ a classical $2$-bit string $r_{1}r_{2}$. There are two qubits per block, initially in the ground state $|00\rangle$. The following state is then obtained, i.e., the corresponding Bell state for the classical bit string.
@@ 4.4 Consensus
[…]
@@ 4.5 Security and Integrity
4.5.1 Theoretical Tamper Sensitivity/Security of the GHZ Blockchain in Comparison to the Quantum Hypergraph Blockchain[2]
It is important to note that the weighted hypergraph states used in [2] do not possess the same degree of multipartite entanglement as maximally entangled GHZ-type states. A phased GHZ state,
$|{\rm GHZ}_{\phi}\rangle=\frac{1}{\sqrt{2}}\left(|0\ldots 0\rangle+e^{i\phi}|1\ldots 1\rangle\right),$
remains maximally entangled for any value of the phase $\phi$, and is extremely fragile under local measurements: a projective measurement on any single qubit collapses the entire state and destroys the entire multipartite entangled state.
## Start of [41]
The "Non"-Maximal Entanglement of Weighted Hypergraph States[41]
According to [41], a weighted graph state (WGS), a subclass of Weighted Quantum Hypergraph States (WQHS), is generated when controlled-phase (CP) gates with an angle $\phi\neq\pi$ are used, instead of the maximally entangling Controlled-Z (CZ) gates ($\phi=\pi$). This implies that the resulting state is generally not maximally entangled. This paper, "Measurement-based quantum computation on weighted graph states with arbitrarily small weight,"[41] explicitly addresses the nature of the entangling gates and the resulting state:
"Weighted graph states are a natural generalization of graph states, which are generated by applying controlled-phase gates, instead of controlled-Z gates, to a separable state. … One such experimental constraint is that only controlled-phase (CP) gates with angle $\phi$ are available, instead of controlled-Z (CZ) gates ($\phi=\pi$). Then, a generated state becomes a so-called **weighted graph state** (WGS), instead of a graph state." [41]
[41] Continuation
They further classify the entanglement inherent to these states, stating that WGS are the first example of a universal resource created this way:
"To our knowledge, this is the first example of universal resources prepared with only **non-maximally entangling gates** and has potential applications to weakly interacting systems, such as photonic systems." [41]
This use of non-maximally entangling gate to prepare the WGS (and, more generally, WQHS with variable weights) suggests that the resulting states may not always be maximally entangled.
## End of [41]
In contrast to GHZ, the weighted hypergraph states employed in the hypergraph blockchain construction [2] have a significantly more distributed and non-uniform entanglement structure. As stated above, these states include amplitudes for all computational basis states,
$|\psi\rangle=\frac{1}{\sqrt{2^{N}}}\sum_{x\in\{0,1\}^{N}}e^{i\pi f(x)}|x\rangle,$
with phases determined by the pattern and weights of multiple $k$-hyperedges. Because the entanglement in such states is typically non-maximal and is spread across many local correlations, their response to local measurements is less catastrophic than that of a GHZ state. In general, measuring a single qubit of a weighted hypergraph state theoretically does not theoretically annihilate all global correlations, and the post-measurement state may still retain nontrivial entanglement among the remaining qubits.
Thus, from a theoretical standpoint, GHZ states are more fragile and more sensitive to tampering or measurement, whereas the weighted hypergraph states used in this protocol possess a resilient but non-maximal entanglement (less entangled) structure.
For instance, the 3-hyperedged, 3-qubit hypergraph(blockchain) from [2] is
$|\psi\rangle_{123}=X_{3}C_{12}^{2}C_{13}^{2}\,|\psi\rangle_{12}\otimes|\psi\rangle_{3}$
$\frac{1}{2\sqrt{2}}\Big(e^{i\theta_{p_{3}}}|000\rangle+|001\rangle+e^{i(\theta_{p_{2}}+\theta_{p_{3}})}|010\rangle+e^{i\theta_{p_{2}}}|011\rangle+e^{i(\theta_{p_{1}}+\theta_{p_{3}})}|100\rangle+e^{i\theta_{p_{1}}}|101\rangle+e^{i(\theta_{p_{1}}+\theta_{p_{2}}+\theta_{p_{3}})}|110\rangle-e^{i(\theta_{p_{1}}+\theta_{p_{2}})}|111\rangle\Big).$
, where it is intuitive that, for any $\theta_{p_{1}},\theta_{p_{2}},\theta_{p_{3}}$ chosen according to the consensus protocol in [2], this state is not as entangled as a GHZ state. It is worth emphasizing that although weighted hypergraph states belong to the class of LME states, this property does not in general imply maximal multipartite entanglement. In the sense used in this work, LME refers to the ability of a state to generate maximal entanglement between the system and a set of local ancillas via suitable local unitary operations. This is an operational notion that concerns the entangling power of the state under local control, rather than the intrinsic amount of global entanglement present in the multipartite system itself. In particular, the degree of genuine multipartite entanglement depends sensitively on the choice of hyperedge weights and the resulting phase structure. For generic weight assignments, the reduced density matrices of subsystems need not be maximally mixed, and standard entanglement measures need not attain their maximal values. Consequently, while the LME property ensures strong nonclassical correlations and local controllability, it does not guarantee that the corresponding weighted hypergraph states are maximally entangled in the multipartite sense.
Hence, a GHZ blockchain is theoretically more tamper/measurement-sensitive and secure than the hypergraph blockchain[2].
@@ 4.5.2 The Spatial GHZ Blockchain Security: Without the Temporal Aspect
The data is encoded into the phase of the GHZ state and the basis states themselves. Therefore, for an adversary to measure the phase, they will need to keep measuring the state in random basis as long as the length of the chain, $\theta_{P_{1}}$, classical 2-bit strings, or $n$ remain secrets. This randomness includes GHZ basis states computational basis states as well. Even if the chain is constructed by the adversary using those parameters (even if he obtains bit strings and the phases of blocks), without the knowledge of the bijective functions, it is not possible to obtain the data. Even if the data is obtained, i.e, the bijective functions are discovered, it is still impossible to mutate the data with that same bijective function intact, i.e, change the phase of the $i$th block - $\theta_{Pi}$ - and the classical string - $r_{i_{1}}r_{i_{2}}$ -, since this would result in a different phase than that which is predefined for the GHZ state and different basis states that don’t represent the classical bit strings $r_{i_{1}}r_{i_{2}}$. If the length of the chain is $m$, i.e. $2m$ qubits, the projection measurement basis to check validity will be
@@
with the rest of the basis set: the other $2^{2n}-2$ states that make an orthonormal basis (GHZ basis with an added phase) with these 2 states, where in the projection $\Omega,I-\Omega$}, $\Omega$ is the projection operator for the ++ state out of the two states above.
@@ 6 Conclusion
7 Limitations
It must be understood that this work is to merely convey a conceptual blockchain on a quantum register. Low-level, case-by-case attacker and security models, a quantum consensus algorithm tailored to the phased GHZ and Bell states in question must be constructed. This work is not intended to be a fully formalized, rigorous theoretical framework or an engineering-ready design of a blockchain.
  1. 1.
    Lack of full consensus protocol formalization. While we outline a high-level consensus mechanism, a fully detailed quantum consensus protocol with rigorous message flows, fault tolerance proofs, and an explicit adversarial model in a quantum network remains to be done. We invite enthusiastic academics to build upon our base. Similar to [1], the starting point for this development could consider the $\theta$-protocol [47].
  2. 2.
    Adversary and threat model constraints. The security analyses assume an idealized quantum network, limited error models, and simplified adversarial capabilities. A comprehensive, case-by-case attacker model (including quantum memory attacks, denial-of-service, network partitioning, and measurement coercion) is not fully treated.
  3. 3.
    Technology and implementation readiness. The proposal leverages time entanglement, phase encoding, and phased GHZ states. No empirical prototype or detailed engineering study on current photonic hardware, error rates, or scalability in practical deployment is included. All features adopted from [1] have been proven in the empirical sense as claimed by [1]. Feasibility of large-scale quantum-register blockchain networks remains to be demonstrated. All features adopted from [2] are empirically proven through their simulations. It remains to be empirically proven whether phase can be added to spatial and temporal GHZ states.
  4. 4.
    Scalability and resource analysis. While combining time-entanglement with phase encoding aims to enhance efficiency and scalability of the temporal GHZ blockchain [1], detailed quantitative resource estimates (e.g., qubit counts, entanglement generation overhead, decoherence thresholds, throughput/latency metrics) for realistic network sizes are not provided by us.
  5. 5.
    Fusion process using PBS In [16], they successfully produced GHZ states by performing Bell projection on photons from different Bell pairs using polarized beam splitting. This was done without considering the phase. Our framework requires strict phase control. This fusion with precise phase control needs to be empirically demonstrated.
  6. 6.
    Classical-quantum interface and operational integration. The transition between classical data and the quantum blockchain register is described conceptually, but a full operational specification (including classical-quantum interaction layers, transaction validation mechanisms, revocation models, and hybrid network protocols) is left for future work.